# Another look at inference after prediction

### v4

@@ 7 Discussion
In this paper, we bridge classical statistical methods with a modern problem to equip practitioners and methodologists with a rigorous approach for PB inference. We find that incorporating a simple weight in the PPI estimator previously proposed for studies with double-sampling designs22 yields valid inference that is guaranteed to be at least as efficient as estimation using only labeled data. Our work addresses efficiency shortcomings of the original PPI estimator which can undermine its practical value. Alongside theoretical results supporting these findings, our simulation studies and application to UK Biobank data verify that the PPI method can suffer substantial efficiency loss when the ML model is inaccurate. Additionally, we draw important connections to several established approaches from the economics and statistics literature to provide context for ongoing developments in PB inference.
Our work is not without limitations. First, we focused our analyses on least squares estimation following Motwani and Witten.20 Further work is needed to analyze other statistical procedures (e.g., Z-estimation problems in general). Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random. This assumption may not hold in practice, particularly when samples cannot be randomly selected for labeling due to study design constraints or the use of alternative sampling strategies to enhance efficiency (e.g., stratified sampling). However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment. Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0. Analyses of semi-supervised settings where $\lim_{n_{\text{\tiny{lab}}}\to\infty}n_{\text{\tiny{lab}}}/n=0$ would be of interest to formally model scenarios where the size of labeled and unlabeled are of different orders of magnitude.
@@
Our analyses can be replicated with the code at: https://github.com/SelinaS37/PBInference.

### v5

@@ 7 Discussion
In this paper, we expand on existing research into the efficiency of PB inference methods by making the following contributions:
  1. 1.
    We employed the AIPW framework to introduce the CC estimator from the double-sampling literature as a simple modification of the PPI estimator, offering theoretically guaranteed efficiency gains over classical inference. Additionally, we demonstrated that the CC estimator is the most efficient estimator within the class of weighted linear augmentation estimators.
  2. 2.
    We made numerous connections between PB inference and existing literature with direct implications for future methodology and applications. Of note, we provided a set of results relating newly proposed PB inference methods to the CC method. For linear regression, we showed that the PPI++ and PSPA estimators are asymptotically equivalent to weighted linear augmentation estimators and therefore the CC estimator is always at least as efficient as these two methods. This theoretical finding was corroborated in our simulated and real data analyses.
  3. 3.
    We conducted the most extensive simulation study to date comparing existing PB inference approaches. We demonstrated empirically that the CC estimator outperforms the PPI++ and PSPA estimators in numerous simulation scenarios as well as in an analysis of UK Biobank data.
However, our work is not without limitations. First, we focused our analyses on Z-estimation and further work is needed to analyze other statistical procedures. Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random. This assumption may not hold in practice, particularly when samples cannot be randomly selected for labeling due to study design constraints or the use of alternative sampling strategies to enhance efficiency (e.g., stratified sampling). However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment. Third, our current analysis assumes that only one prediction is available for analysis. In many practical scenarios multiple predictions with unknown accuracy are available to the analyst. The CC estimator can be naturally extended to accommodate multiple predictions and we provide a simulated example in the Appendix. Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0. Analyses of semi-supervised settings where $\lim_{n_{\text{\tiny{lab}}}\to\infty}n_{\text{\tiny{lab}}}/n=0$ would be of interest to formally model scenarios where the size of labeled and unlabeled are of different orders of magnitude.
@@
Our analyses can be replicated with the code at: https://github.com/jlgrons/PBInf-Simulations.